### v1

@@ Abstract
This paper proposes a probabilistic machine learning method to price catastrophe (CAT) bonds in the primary market. The proposed method combines machine-learning-based predictive models with Conformal Prediction, an innovative algorithm that generates distribution-free probabilistic forecasts for CAT bond prices. Using primary market CAT bond transaction records between January 1999 and March 2021, the proposed method is found to be more robust and yields more accurate predictions of the bond spreads than traditional regression-based methods. Furthermore, the proposed method generates more informative prediction intervals than linear regression and identifies important nonlinear relationships between various risk factors and bond spreads, suggesting that linear regressions could misestimate the bond spreads. Overall, this paper demonstrates the potential of machine learning methods in improving the pricing of CAT bonds.
Keywords and phrases: Catastrophe Bond, Machine Learning, Conformal Prediction, Probabilistic Forecasting, Nonlinear Pricing
@@ 1 Introduction
Catastrophe (CAT) bonds have become increasingly vital in managing and transferring catastrophic risk. These bonds offer a source of capital to cover losses arising from natural disasters, allowing investors to diversify their portfolios while helping issuers mitigate potentially devastating financial consequences. Understanding the pricing dynamics of CAT bonds is essential, both for investors seeking informed decisions and for issuers optimizing their risk management strategies. This paper introduces a probabilistic machine-learning-based predictive framework for the pricing of CAT bonds, aiming to enhance empirical pricing accuracy and discover previously undetected nonlinear dependence between the key risk factors and CAT bond spreads.
@@
Early research by Lane, (2000) laid the groundwork for CAT bond pricing literature, proposing a log-linear regression model employing conditional expected loss and probability of first loss as predictors. Subsequent studies expanded on this linear framework, incorporating additional predictors and examining pricing under diverse conditions. Gürtler et al., (2016) incorporated bond characteristics like trigger type and bond rating, while Braun, (2016) integrated market condition indices, such as the Lane Synthetic Rate on Line index and the BB corporate bond spread. Götze and Gürtler, 2020a explored sponsor-related pricing inefficiencies across different market conditions, and Morana and Sbrana, (2019) focused on the impact of climate change on CAT bond returns. Further extending the research scope, Zhao and Yu, (2020) utilized actual catastrophe data to forecast CAT bond prices using market-based methods, Braun et al., (2022) developed factor pricing models for cross-sectional CAT bond returns, and Herrmann and Hibbeln, (2023) investigated liquidity premiums in the secondary market. Chatoro et al., (2023) explored issuer effects in the primary market. However, relying on linear models may lead to oversimplifications, disregarding nonlinear relationships and predictor interactions, potentially limiting predictive accuracy.
To overcome this limitation, recent literature has ventured into more flexible predictive methods. One line of research has focused on the development of structural models for CAT bond pricing. For example, Lee and Yu, (2002) studied the pricing of default-risky CAT bonds in the presence of moral hazard and basis risk; Zhu, (2017, 2011) proposed pricing formulas by allowing agents to act in a robust control framework against model misspecification; Lee and Yu, (2007) proposed a contingent-claim framework for valuing catastrophe reinsurance contracts and catastrophe bonds.
@@
More recently, machine learning models have been introduced to CAT bond pricing, aiming to better capture the nonlinear dynamics and interactions between predictors. Makariou et al., (2021) introduced the Random Forest method to predict CAT bond spreads which produced superior predictive performance compared to traditional linear models. In addition, Götze and Gürtler, 2020b and Götze et al., (2023) applied Random Forest and Neural Network techniques to improve the forecasting accuracy of CAT bond prices.
@@
Machine learning methods often focus on point forecasts. However, considering that predictions often deviate from the actual prices due to market complexity and dynamism, exploring both points forecasts and prediction intervals is equally crucial. Generating prediction intervals for asset prices is standard practice in financial markets. When issuing a CAT bond, the process typically involves an initial offering where the issuer provides a range of spreads to potential investors. Probabilistic forecasts can help the issuer gauge how likely the final spread will fall within a specific range. Repricing can also happen during the issuing process based on investor feedback and market conditions, thereby highlighting the inadequacy of relying solely on point forecasts for informed decision-making and effective risk management. Bond issuers, who seek to transfer risks, should consider not just a single estimate but a credible range of potential costs, including bond spreads in unfavorable scenarios, to realistically appraise the cost of risk transfer. For investors, understanding the possible range where the actual price could fall within is also vital in making informed investment decisions.
While linear regression models readily facilitate the generation of prediction intervals with underlying distributional assumptions, machine learning techniques like Random Forest and Neural Network, while providing more precise price estimates, typically do not yield prediction intervals. This paper aims to combine the accuracy of machine learning predictions with the benefit of probabilistic forecasts in CAT bond pricing.
@@
Our first methodological contribution introduces the use of Extreme Gradient Boosting (XGBoost) (Chen and Guestrin,, 2016) for predicting CAT bond spreads. XGBoost, a robust machine learning tool, employs sequential boosting of weak learners (typically decision trees) to create a strong predictive model. While well-established in other domains, XGBoost has not been applied to CAT bond pricing. Our research explores its potential by comparing its prediction accuracy with that of the widely used linear regression pricing models and Random Forest. We also employ various interpretation tools, including feature importance and accumulated local effects plots, to understand the nonlinear relationships and complex interactions between different risk factors and bond spreads.
@@
The second major contribution of our study applies Conformal Prediction for constructing prediction intervals around point forecasts from machine learning models. Conformal Prediction intervals, studied by Shafer and Vovk, (2008), Angelopoulos and Bates, (2021), and Fontana et al., (2023), are constructed from the residuals of the trained machine learning model and offer guaranteed coverage of the true response values with user-specified probability. Conformal Prediction is distribution-free and highly flexible, applicable to any prediction model, including Random Forest, Gradient Boosting algorithms, and various Neural Networks. It can even be applied to linear regression models, although this offers little value since producing prediction intervals is straightforward for linear regression models.
@@
Empirically, our study analyzes primary CAT bond market transaction records from January 1999 to March 2021, utilizing data from reputable sources including trade notes provided by Lane Financial LLC, the Artemis Deal Directory, reports from rating agencies, and market research conducted by prominent industry participants such as Swiss Re, Munich Re, Aon Benfield, and Guy Carpenter. Similar data sets with smaller sample sizes were used in existing studies, such as Braun, (2016). Our results highlight several advantages of the XGBoost model over traditional linear regression models. Firstly, XGBoost consistently achieves higher out-of-sample forecasting accuracy and generates more informative prediction intervals through Conformal Prediction. These prediction intervals are, on average, 8% narrower than those from linear models, yet they maintain comparable coverage rates. Furthermore, XGBoost effectively handles predictors with high correlations, such as expected loss and probability of first loss, while linear regression models often yield unstable estimates and interpretational challenges due to highly correlated predictors. In linear regression, the usual practice involves removing one of the correlated predictors, but this approach could inadvertently eliminate valuable information inherent in the excluded variable, leading to a less optimal model. XGBoost preserves the information they carry and enhances the model performance by allowing for highly correlated variables. Lastly, our approach reveals the nonlinear effects of several important predictors on CAT bond spreads, including expected loss, probability of first loss, issue size, and reinsurance market cycles. We also observe significant interactions between these predictors. For example, we find that the spread of bonds with higher expected loss values is more sensitive to reinsurance market conditions. In addition, the impact of issue size is negatively correlated with expected loss for bonds with high expected loss values. These nonlinear relationships and complex interactions are naturally captured by the XGBoost model. In summary, this paper extends the existing literature on CAT bond pricing by introducing a probabilistic machine learning approach that improves forecasting accuracy, and provides probabilistic forecasts as well as insights on the nonlinear effects of various risk factors on CAT bond spread.
@@
The structure of this paper is as follows: Section 2 outlines the methodologies employed in our study, including linear regression, ensemble learning techniques (Random Forest and XGBoost), and the Conformal Prediction approach. Section 3 provides an overview of the CAT bond transaction data and an exploratory analysis. Section 4 presents empirical analysis results and their interpretations. Section 5 constructs Conformal Prediction intervals and conducts the probabilistic analysis of CAT bond spreads. Finally, Section 6 concludes.
2 Methodology
This section introduces the predictive methods used in our analysis. Section 2.1 formulates the bond-yield prediction problem, while Sections 2.2 and 2.3 introduce the linear model and ensemble machine learning methods, respectively. Finally, Section 2.4 outlines the Conformal Prediction approach.
2.1 Problem Formulation
Let us denote the spread of a CAT bond at issuance as $Y$, and let $\bm{X}=(X_{1},X_{2},\ldots,X_{p})$ represent $p$ distinct predictors. This predictor vector $\bm{X}$ encompasses variables that convey information about bond characteristics, market conditions, and economic indicators. We assume there exists a relationship between $Y$ and $\bm{X}$, expressed as:
$Y=f(\bm{X})+\epsilon.$
Assume that we have observed a collection of $n$ distinct data points from primary CAT bond market transactions. We use $x_{ij}$ to represent the value of the $j$th predictor from the $i$th observation, where $i=1,2,\ldots,n$ and $j=1,2,\ldots,p$. Correspondingly, $y_{i}$ denotes the value of the response variable from the $i$th observation. We can represent the observed dataset as $D=\{(\bm{x}_{1},y_{1}),(\bm{x}_{2},y_{2}),\ldots,(\bm{x}_{n},y_{n})\}$, where $\bm{x}_{i}=(x_{i1},x_{i2},\ldots,x_{ip})^{T}$.
The estimation of $f$ involves exploring the hypothesis space $H$, which is the set of all feasible functions, to find a function, denoted as $\hat{f}$, that can predict the response for a new input as accurately as possible. This process, commonly referred to as learning, is guided by minimizing a loss function that quantifies the difference between the true output $(y_{1},y_{2},\ldots,y_{n})$ from the observed data and the predicted outputs $\bigl{(}\hat{f}(\bm{x}_{1}),\hat{f}(\bm{x}_{2}),\ldots,\hat{f}(\bm{x}_{n})\bigr{)}$.
2.2 Linear Regression
$f(\bm{X})=\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+\cdots+\beta_{p}X_{p}.$
One advantage of linear regression is its simplicity and ease of interpretation. However, this assumed linear relationship may not always accurately represent the true underlying relationship, which could lead to substantial errors in the predictions.
2.3 Ensemble Learning
This subsection introduces two ensemble learning methods used in our analysis: Random Forest and XGBoost. An overview of ensemble methods is provided in Section 2.3.1, while Sections 2.3.2 and 2.3.3 explore the specifics of the two ensemble methods.
2.3.1 Overview
Ensemble learning is a widely adopted machine learning technique that leverages the power of multiple models (base learners) to collectively address a problem, resulting in improved accuracy and robustness compared to individual models. Instead of relying on a single function $f$ for estimation, ensemble learning combines the outputs of multiple functions that approximate $f$. In this study, we employ two well-established ensemble learning methods: Random Forest and XGBoost.
Ensemble learning uses base learners, often capable of capturing nonlinearity, such as decision trees, to model complex relationships between predictors and the response variable. Consequently, ensemble learning frequently outperforms linear regression in prediction accuracy. However, the amalgamation of multiple models can make the output less interpretable. To better understand the model’s behavior, various strategies, including partial dependence plots and feature importance measures, have been developed to illustrate the significance of individual predictors in model predictions. Nevertheless, the interpretability of an ensemble model remains less straightforward compared to that of a single decision tree or linear regression model.
2.3.2 Random Forest
The Random Forest algorithm, introduced by Breiman (Breiman,, 2001), is a potent machine learning technique that combines two strategies: bagging (Breiman,, 1996) and random feature selection (Ho,, 1998, 1995; Amit and Geman,, 1997). The construction of a Random Forest model comprises three steps. First, it generates $K$ bootstrapped samples from the original dataset $D$. Next, it constructs a decision tree for each bootstrapped sample, using recursive binary splitting with random feature selection. This results in $K$ estimators of the function $f$, denoted as $\hat{f}_{1},\hat{f}_{2},\ldots,\hat{f}_{K}$. Finally, to create the Random Forest estimator, these individual estimators are averaged, yielding $\hat{f}\left(\bm{X}\right)=\frac{1}{K}\sum_{k=1}^{K}\hat{f}_{k}\left(\bm{X}\right)$.
Random Forest typically employs fully grown and unpruned decision trees. Averaging results from multiple trees mitigates overfitting, which is a common issue with fully grown trees. Including more trees in a Random Forest model provides a more diverse set of predictions. This, in turn, reduces variance, resulting in a more robust and accurate model. However, the improvement in performance plateaus after a certain number of trees, where additional computation does not lead to substantial gains in model performance. Thus, the number of trees should be carefully chosen to balance computational efficiency and model performance. To increase the difference among trees and thus mitigate overfitting, Random Forest employs column sub-sampling, meaning that it selects only a subset of predictors for each tree.
2.3.3 XGBoost
2.4 Conformal Prediction
2.4.1 Prediction Interval
$P\bigl{(}y_{n+1}\in C_{\alpha}(\bm{x}_{n+1})\bigr{)}\geq 1-\alpha.$ (2.1)
2.4.2 Overview of Conformal Prediction
$P(u_{n+1}\leq\hat{q}_{1-\alpha}\{u_{1},\ldots,u_{n}\})\geq 1-\alpha,$
where $u_{(1)}\leq\ldots\leq u_{(n)}$ denote the order statistics of $u_{1},\ldots,u_{n}$, and $\lceil(n+1)(1-\alpha)\rceil$ is the smallest integer larger or equal to $(n+1)(1-\alpha)$.
2.4.3 Jackknife+ Conformal Prediction Interval
  1. 1.
  2. 2.
  3. 3.
    Repeat Steps 1 and 2 for all $i=1,\ldots,n$.
  4. 4.
  5. 5.
  6. 6.
While the Jackknife method often empirically achieves $1-\alpha$ coverage, it does not guarantee finite sample coverage validity. Asymptotic validity is only ensured when non-trivial conditions are satisfied by the trained model (Steinberger and Leeb,, 2016). The Jackknife prediction interval may exhibit poor coverage when the trained model $\hat{f}$ is unstable due to its lack of finite sample coverage validity.
@@ 3 Catastrophe Bond Data
Our empirical study focuses on the primary market transactions of CAT bonds. A significant challenge in empirical research on CAT bond pricing is the limited availability of transaction data. To tackle this issue, we adopted the strategy outlined by Braun, (2016), which involves aggregating and verifying information from diverse sources. The dataset of CAT bonds used in this study encompasses a total of 765 bond tranches issued between January 1999 and March 2021. This dataset includes various critical bond characteristics, such as expected loss, probability of first loss, covered perils and territories, the trigger mechanism, sponsor information, and credit rating. The data was compiled from trade notes sourced from Lane Financial LLC [^1] Source: http://lanefinancialllc.com/., the Artemis Deal Directory [^2] Source: https://www.artemis.bm/deal-directory/., reports from credit rating agencies, and market research conducted by Swiss Re, Munich Re, Aon Benfield, and Guy Carpenter.
We begin with an exploratory analysis of the dataset. Figures 1(a) and 1(b) depict scatter plots of bond spread in relation to expected loss and the probability of first loss. We can see a strong positive correlation between bond spread and both expected loss and probability of first loss in the sample. While the majority of bonds exhibit expected loss values below 5% and probabilities of first loss below 10%, there are some bonds with relatively high values of these factors. Furthermore, a handful of bonds demonstrate exceptionally large spreads relative to their expected loss or probability of first loss. This observation reveals that factors beyond expected loss and probability of first loss may affect the bond spread.
@@
Figure 2(a) offers an overview of the volume and number of new transactions per year in the primary CAT bond market. We observe a sharp drop in CAT bond volume in 2008 during the financial crisis. This decline can be attributed to the flight-to-quality phenomenon, where investors tend to avoid risky assets during periods of market stress (Caballero and Krishnamurthy,, 2008). However, investor confidence gradually rebounded, leading to a resurgence in CAT bond issuance in 2009 and 2010, with peak activity occurring in 2020. The surge in CAT bond issuance in 2020 was largely influenced by the COVID-19 pandemic, which motivated new issuers to enter the CAT bond market (Chatoro et al.,, 2023). Figure 2(b) presents the total issuance volume and the number of new transactions per calendar month. This chart demonstrates that the primary CAT bond market experiences heightened activity during the first, second, and fourth quarters of the year, while activity tends to be quieter in the third quarter. As argued by Braun, (2016), the upsurge in issuance volumes towards the end of the calendar year may be linked to sponsors using securitization to rebalance their firm-wide risk positions before preparing their financial statements. It is interesting to note that our findings differ from those of Braun, (2016), which suggested relative inactivity in the first quarter. This discrepancy may arise from the inclusion of more recent transaction data in our study, which extends up to 2021, compared to the data used in Braun, (2016)’s analysis, which covers up to December 2012. Over the past decade, the frequency and severity of natural disasters have increased, potentially contributing to greater demand for coverage in the first quarter. The expansion of the insurance-linked securities market is another potential contributor to the increased activity in the first quarter.
@@
In addition to bond-specific data, we consider several market-level indices to proxy the reinsurance market and interest rate cycles. We employ the Lane Financial LLC Synthetic Rate on Line Index ($\mathrm{ROLX}$), published quarterly by Thomson Reuters ILS Community, to represent the global reinsurance cycle. $\mathrm{ROLX}$ integrates data from the catastrophe bond, insurance-linked security (ILS), and industry-loss warranty (ILW) markets, thereby reflecting the average premiums paid for ILS and CAT bond-backed reinsurance or retrocession. In addition, we utilize the Guy Carpenter (GC) Global Property Catastrophe Rate on Line indices, covering different regions—Global, US, Asia Pacific, Europe, and UK—as indicators of the premium levels for reinsurance specific to catastrophe risk. These GC indices, published annually, monitor property catastrophe reinsurance rate-on-line movements in major global catastrophe reinsurance markets. They are based on reinsurance costs and reflect the expenses incurred in protecting property against catastrophic losses. The GC indices were previously employed by Gürtler et al., (2016) as predictors of CAT bond spreads. A larger value of these reinsurance market indices suggests an upswing in the reinsurance market cycle (i.e., a hard market), when premiums increase, coverage terms are restricted, and capacity for most types of reinsurance decreases. Furthermore, we incorporate the Bank of America Merrill Lynch US High Yield BB Option-Adjusted Spread (BB spread), which is published daily, to capture the influence of the corporate bond markets. This index is calculated as the difference between a yield index for the BB rating category and the Treasury spot curve.
@@
Tables 3.1 and 3.2 present a range of descriptive statistics for our sample of 765 CAT bonds. Table 3.1 displays the mean, median, standard deviation, maximum, and minimum values for some key variables in the sample. The mean spread across all bonds is 771.38 bps, and the mean expected loss is 244.01 bps. The average risk premium per bond is 527.36 bps, calculated as the difference between spread and expected loss. The spread exhibits considerable variability with values ranging from 66 bps to 4,988 bps. The highest spread is reported by Successor Hurricane Industry Ltd. (Series 3), which was identified as an outlier in Braun, (2016). However, since our analysis focuses on machine learning methods that are more robust to outliers than linear regression models, we choose to retain this record in our sample. Similarly, expected losses vary widely, spanning from 1 bps to 1,484 bps. The sample includes CAT bond tranches with sizes ranging from USD \$1.80 million to USD \$1.5 billion, with a median size of USD \$100 million. The CAT bond tranche with an exceptionally large size of \$1.5 billion is from Everglades Re 2014-1 A which covers Florida hurricanes. Terms for CAT bonds in the sample vary from 5 months to 73 months, with a median maturity period of 37 months.

### v2

@@ Abstract
This paper explores the implications of using machine learning models in the pricing of catastrophe (CAT) bonds. By integrating advanced machine learning techniques, our approach uncovers nonlinear relationships and complex interactions between key risk factors and CAT bond spreads—dynamics that are often overlooked by traditional linear regression models. Using primary market CAT bond transaction records between January 1999 and March 2021, our findings demonstrate that machine learning models not only enhance the accuracy of CAT bond pricing but also provide a deeper understanding of how various risk factors interact and influence bond prices in a nonlinear way. These findings suggest that investors and issuers can benefit from incorporating machine learning to better capture the intricate interplay between risk factors when pricing CAT bonds. The results also highlight the potential for machine learning models to refine our understanding of asset pricing in markets characterized by complex risk structures.
Keywords and phrases: Catastrophe Bond, Asset Pricing, Machine Learning, Nonlinear Relationships, Conformal Prediction
@@ 1 Introduction
Catastrophe (CAT) bonds play a crucial role in transferring and managing the financial risks associated with natural disasters. These instruments provide a valuable source of capital for issuers, helping them mitigate potentially devastating financial losses, while offering investors an opportunity to diversify their portfolios. Understanding how CAT bond prices are determined hence is essential for both investors, who seek to make informed decisions, and issuers, who aim to optimize their risk management strategies. However, accurately pricing CAT bonds is a challenging task, as prices are influenced by a wide array of factors, including bond-specific characteristics and prevailing market conditions. This paper adopts a machine-learning-based approach to explore the complex relationships driving CAT bond prices, with a focus on uncovering potential nonlinear impacts and interaction effects among key risk factors. By shedding light on these intricate dynamics, the study contributes to a deeper understanding of asset pricing in the context of CAT bonds.
@@
Early research by Lane, (2000), which introduced a log-linear regression model, laid the foundation for the CAT bond pricing literature. Lane identified two key determinants of CAT bond prices in the primary market: the probability of first loss—the likelihood that the bond will experience any loss of principal due to a triggering event—and the conditional expected loss, which is the average percentage of the bond principal expected to be lost if the bond is triggered. Subsequent studies built on this linear framework by incorporating additional predictors and exploring pricing under various conditions. For instance, Gürtler et al., (2016) included bond characteristics such as trigger type and bond rating, while Braun, (2016) integrated market condition indices including the Lane Synthetic Rate on Line index and the BB corporate bond spread. Götze and Gürtler, 2020a investigated sponsor-related pricing inefficiencies across different market conditions, and Morana and Sbrana, (2019) examined the impact of climate change on CAT bond returns. Expanding the research scope further, Braun et al., (2022) developed factor pricing models for cross-sectional CAT bond returns, and Herrmann and Hibbeln, (2023) analyzed liquidity premiums in the secondary market. Chatoro et al., (2023) also explored issuer effects in the primary market. However, relying solely on linear models can oversimplify CAT bond price dynamics by overlooking nonlinear relationships and interactions between predictors.
In addition to using risk factors to predict the price of CAT risks, several other approaches have been proposed. For instance, Lee and Yu, (2002) examined the pricing of default-risky CAT bonds in the context of moral hazard and basis risk; Niehaus, (2002) explored the efficiency (or lack thereof) in the allocation of catastrophe risk among individuals, primary insurers, and reinsurers; Duan and Yu, (2005) analyzed how catastrophe risk affects the fair premiums for insurance guaranty funds; Froot and O’Connell, (2008) investigated the role of intermediated risks in catastrophe reinsurance; Chang et al., (2010) priced an Asian CAT option using a non-traded underlying loss index; Lee and Yu, (2007) introduced a contingent-claim framework for valuing catastrophe reinsurance contracts and CAT bonds; Perrakis and Boloorforoosh, (2013) proposed a stochastic dominance approach to price catastrophe derivatives under limited diversification; Braun et al., (2019) examined the determining factors affecting the returns of funds specializing in insurance-linked securities; Zhao and Yu, (2020) used actual catastrophe data to forecast CAT bond prices through market-based methods; and Beer and Braun, (2022) estimated smooth intensity surfaces for catastrophe arrivals from secondary CAT bond data.
@@
More recently, machine learning techniques have been widely adopted in the financial asset pricing literature due to their ability to capture complex patterns and relationships (see, among many others, Dong et al.,, 2022; Jiang et al.,, 2023; Kelly et al.,, 2024). This trend is also evident in CAT bond pricing, where machine learning models have been increasingly utilized to better capture the nonlinear dynamics and interactions between predictors. For example, Makariou et al., (2021) introduced the Random Forest method to predict CAT bond spreads, demonstrating superior predictive performance compared to traditional linear models. Additionally, Götze and Gürtler, 2020b and Götze et al., (2023) applied Random Forest and Neural Network techniques to enhance the goodness-of-fit of CAT bond prices.
@@
While machine learning methods could cope with pricing dynamics beyond linear relationships, they often focus on point forecasts. However, considering that predictions often deviate from the actual prices due to market complexity and dynamism, incorporating uncertainties surrounding points forecasts is crucial. Generating prediction intervals for asset prices is standard practice in financial markets. When issuing a CAT bond, the process typically involves an initial offering where the issuer provides tentative spreads to potential investors. Repricing can also happen during the issuing process based on investor feedback and market conditions. Therefore, bond issuers should consider not just a single estimate but a credible range of potential final spreads, especially spreads in unfavorable scenarios, to realistically appraise the cost of risk transfer. For investors, understanding the possible range where the actual price could fall within is also vital in making informed investment decisions. This paper aims to combine the superior ability of machine learning approaches to discover complex pricing dynamics with the benefit of probabilistic forecasts in CAT bond pricing.
@@
Our first methodological contribution introduces the use of Extreme Gradient Boosting (XGBoost) (Chen and Guestrin,, 2016) for modeling CAT bond spreads. XGBoost, a robust machine learning tool, employs sequential boosting of weak learners (typically decision trees) to create a strong predictive model. While well-established in other domains, XGBoost has not been applied to CAT bond pricing. Our research explores its potential by comparing its prediction accuracy with that of the widely used linear regression pricing models and Random Forest. More importantly, we also employ various interpretation tools, including feature importance and accumulated local effects plots, to understand the nonlinear relationships and complex interactions between different risk factors and bond spreads.
@@
The second contribution of our study applies Conformal Prediction for constructing prediction intervals around point forecasts from machine learning models. Conformal Prediction intervals, studied by Shafer and Vovk, (2008), Angelopoulos and Bates, (2021), and Fontana et al., (2023), are constructed from the residuals of the trained machine learning model and offer guaranteed coverage of the true response values with user-specified probability. Conformal Prediction is distribution-free and highly flexible, applicable to any prediction model, including Random Forest, Gradient Boosting algorithms, and various Neural Networks. It can even be applied to linear regression models, although this offers little value since producing prediction intervals is straightforward for linear regression models.
@@
Empirically, our study analyzes primary CAT bond market transaction records from January 1999 to March 2021, utilizing data from multiple data sources including trade notes provided by Lane Financial LLC, the Artemis Deal Directory, reports from rating agencies, and market research conducted by prominent industry participants such as Swiss Re, Munich Re, Aon Benfield, and Guy Carpenter. Our results highlight several advantages of the XGBoost model over traditional linear regression models. First, our approach reveals the nonlinear effects of several important risk factors on CAT bond spreads, including expected loss, probability of first loss, issue size, and reinsurance market cycles. We also observe significant interactions between these risk factors. For example, we find that the spread of bonds with higher expected loss values is more sensitive to reinsurance market conditions. In addition, the impact of issue size is negatively correlated with expected loss for bonds with high expected loss values. These nonlinear relationships and complex interactions are disregarded by linear regression models but naturally captured by the XGBoost model. Furthermore, XGBoost effectively handles risk factors with high correlations, such as expected loss and probability of first loss, while linear regression models often yield unstable estimates and interpretational challenges due to highly correlated risk factors. In linear regression, the usual practice involves removing one of the correlated risk factors, but this approach could inadvertently eliminate valuable information inherent in the excluded variable, leading to a less optimal model. XGBoost preserves the information they carry and enhances the model performance by allowing for highly correlated variables. Last but not least, XGBoost consistently achieves higher out-of-sample forecasting accuracy and generates more informative prediction intervals through Conformal Prediction. These prediction intervals are, on average, 8% narrower than those from linear models, yet they maintain comparable coverage rates. In summary, this paper extends the existing literature on CAT bond pricing by introducing a probabilistic machine learning approach that provides insights on the nonlinear effects of various risk factors on CAT bond spread, as well as probabilistic forecasts with improved accuracy.
@@
The structure of this paper is as follows: Section 2 provides an overview of the primary market CAT bond transaction data and present an exploratory analysis. Section 3 outlines the methodologies employed in our study, including linear regression, XGBoost, and the Conformal Prediction approach. Sections 4 and 5 present the empirical analysis results and their interpretations using linear regression models and machine learning models, respectively. Section 6 constructs Conformal Prediction intervals, evaluates the prediction performance of the two modelling approaches, and further analyse conditional impacts of various risk factors. Finally, Section 7 offers concluding remarks.
@@ 2 Catastrophe Bond Data
The overall return on a CAT bond for the investor consists of two components. First, the investor receives a principal repayment at maturity. If the bond is not triggered, meaning no specified catastrophic events occur during its term, the full principal is repaid. However, if the bond is triggered by one or more specified events, the repayment amount is reduced.
Second, the investor receives coupon payments, which are pre-determined and remain unaffected by whether the bond is triggered. CAT bonds are typically issued as floating-rate securities, where the investor earns a fixed coupon spread over a reference index or the return on high-quality collateral, usually invested in short-term money market funds. The index or collateral return compensates investors for the time value of money and is independent of the bond’s embedded insurance risk. This rate resets periodically based on current short-term interest rates. The spread on a CAT bond compensates investors for assuming the insurance risk associated with potential catastrophic events.
The coupon rate is determined by the following formula:
[…]
Here, the Risk-Free Rate refers to the return on the reference index, while the CAT Bond Spread represents the additional yield that reflects the risk inherent in the bond. A higher CAT bond spread increases the coupon rate, thereby enhancing the overall return to the investor as compensation for the increased risk. Consequently, understanding the determinants of the CAT bond spread is essential for accurate pricing, a topic that has been extensively studied in the literature.
Our empirical study focuses on the primary market transactions of CAT bonds. A significant challenge in empirical research on CAT bond pricing is the limited availability of transaction data. To tackle this issue, we follow existing empirical studies by aggregating and verifying information from diverse sources. The dataset of CAT bonds used in this study encompasses a total of 765 bond tranches issued between January 1999 and March 2021. This dataset includes various critical bond characteristics, such as expected loss, probability of first loss, covered perils and territories, the trigger mechanism, sponsor information, and credit rating. The data was compiled from trade notes sourced from Lane Financial LLC [^1] Source: http://lanefinancialllc.com/., the Artemis Deal Directory [^2] Source: https://www.artemis.bm/deal-directory/., reports from credit rating agencies, and market research conducted by Swiss Re, Munich Re, Aon Benfield, and Guy Carpenter.
We begin with an exploratory analysis of the dataset. Figures 1(a) and 1(b) depict scatter plots of bond spread in relation to expected loss, the average percentage of the bond principal that is anticipated to be lost over the life of the bond due to triggering catastrophic events, and the probability of first loss. We can see a strong positive correlation between bond spread and both expected loss and probability of first loss in the sample. While the majority of bonds exhibit expected loss values below 5% and probabilities of first loss below 10%, there are some bonds with relatively high values of these factors. Furthermore, a handful of bonds demonstrate exceptionally large spreads relative to their expected loss or probability of first loss. This observation reveals that factors beyond expected loss and probability of first loss may affect the bond spread.
@@
Figure 2(a) offers an overview of the volume and number of new transactions per year in the primary CAT bond market. We observe a sharp drop in CAT bond volume in 2008 during the financial crisis. This decline can be attributed to the flight-to-quality phenomenon, where investors tend to avoid risky assets during periods of market stress (Caballero and Krishnamurthy,, 2008). However, investor confidence gradually rebounded, leading to a resurgence in CAT bond issuance in 2009 and 2010, with peak activity occurring in 2020. The surge in CAT bond issuance in 2020 was largely influenced by the COVID-19 pandemic, which motivated new issuers to enter the CAT bond market (Chatoro et al.,, 2023). Figure 2(b) presents the total issuance volume and the number of new transactions per calendar month. This chart demonstrates that the primary CAT bond market experiences heightened activity during the first, second, and fourth quarters of the year, while activity tends to be quieter in the third quarter. It is interesting to note that our findings differ from those of Braun, (2016), which suggested relative inactivity in the first quarter. This discrepancy may arise from the inclusion of more recent transaction data in our study, which extends up to 2021, compared to the data used in Braun’s analysis, which covers up to December 2012. Over the past decade, the frequency and severity of natural disasters have increased, potentially contributing to greater demand for coverage in the first quarter. The expansion of the insurance-linked securities market is another potential contributor to the increased activity in the first quarter.
@@
In addition to bond-specific data, we consider several market-level indices to proxy the reinsurance market and interest rate cycles. We employ the Lane Financial LLC Synthetic Rate on Line Index ($\mathrm{ROLX}$), published quarterly by Thomson Reuters ILS Community, to represent the global reinsurance cycle. $\mathrm{ROLX}$ integrates data from the catastrophe bond, insurance-linked security (ILS), and industry-loss warranty (ILW) markets, thereby reflecting the average premiums paid for ILS and CAT bond-backed reinsurance or retrocession. In addition, we utilize the Guy Carpenter (GC) Global Property Catastrophe Rate on Line indices, covering different regions—Global, US, Asia Pacific, Europe, and UK—as indicators of reinsurance premium levels specific to catastrophe risk. These GC indices, published annually, monitor property catastrophe reinsurance rate-on-line movements in major global catastrophe reinsurance markets. They are based on reinsurance costs and reflect the expenses incurred in protecting property against catastrophic losses. The GC indices were previously employed by Gürtler et al., (2016) as predictors of CAT bond spreads. A larger value of these reinsurance market indices suggests an upswing in the reinsurance market cycle (i.e., a hard market), when premiums increase, coverage terms are restricted, and capacity for most types of reinsurance decreases. Furthermore, we incorporate the Bank of America Merrill Lynch US High Yield BB Option-Adjusted Spread (BB spread), which is published daily, to capture the influence of the corporate bond markets. This index is calculated as the difference between a yield index for the BB rating category and the Treasury spot curve.
@@
Tables 2.1 and 2.2 present a range of descriptive statistics for our sample of 765 CAT bonds. Table 2.1 displays the mean, median, standard deviation, maximum, and minimum values for some key variables in the sample. The mean spread across all bonds is 771.38 bps, and the mean expected loss is 244.01 bps. The average risk premium per bond is 527.36 bps, calculated as the difference between spread and expected loss. The spread exhibits considerable variability with values ranging from 66 bps to 4,988 bps. The highest spread is reported by Successor Hurricane Industry Ltd. (Series 3), which was identified as an outlier in Braun, (2016). However, since our analysis focuses on machine learning methods that are more robust to outliers than linear regression models, we choose to retain this record in our sample. Similarly, expected losses vary widely, spanning from 1 bps to 1,484 bps. The sample includes CAT bond tranches with sizes ranging from USD \$1.80 million to USD \$1.5 billion, with a median size of USD \$100 million. The CAT bond tranche with an exceptionally large size of \$1.5 billion is from Everglades Re 2014-1 A which covers Florida hurricanes. Terms for CAT bonds in the sample vary from 5 months to 73 months, with a median maturity period of 37 months.